### 20 Apr 2026 v3

# Semantic Interaction Information mediates compositional generalization in latent space

Abstract
Are there still barriers to generalization once all of the relevant variables are known? We address this question via a framework that casts compositional generalization as a variational inference problem over latent variables with parametric interactions. To explore this framework, we develop the Cognitive Gridworld, a stationary Partially Observable Markov Decision Process (POMDP) in which observations are generated jointly by multiple latent variables, yet feedback is provided only for a single goal variable. This setting allows us to describe Semantic Interaction Information: an information-theoretic quantity that measures the contribution of latent variable interactions to task performance. Using this metric, we first analyze Recurrent Neural Networks (RNNs) that are explicitly provided with the interactions and find that Semantic Interaction Information explains the accuracy gap between Echo State and Fully Trained networks. Additionally, our analysis uncovers a theoretically predicted failure mode, where confidence becomes decoupled from accuracy. These results suggest that utilizing the interactions between relevant variables is a non-trivial capability.
We then address a harder regime where the interactions themselves must be learned by an embedding model. Learning how latent variables interact requires accurate inference, yet accurate inference depends on knowing those interactions; this circular dependence is characteristic of variational inference. The Cognitive Gridworld reveals this circularity as a potential challenge for continual meta-learning. We approach this dilemma via Representation Classification Chains (RCCs), a JEPA-style architecture which disentangles two processes: variable inference and variable embeddings are learned by separate modules through Reinforcement Learning and self-supervised learning, respectively. Lastly, we demonstrate that, by learning how variables interact, RCCs facilitate compositional generalization to novel combinations of relevant variables and offline learning in novel action spaces. Together, these results establish a theoretically grounded setting for researching, developing and evaluating goal-directed generalist agents.
1 Introduction
As Machine Learning tackles increasingly difficult challenges, rigidity and unreliability remain a significant barrier to the development of generalist agents. A central obstacle is the fact that what can be learned depends crucially on what has been learned. For instance, inferring time of day from ambient light becomes significantly more reliable after learning about seasons (perhaps in order to explain changes in temperature). While a combinatorial explosion can emerge from interactions, only a small number of variables may be relevant to a given task. Therefore, the world could be compressed into something far more manageable, insofar as the goal-directed agent is concerned [100, 81]. Similarly in nature, the knowledge an animal gains from an experience is shaped by internal objectives (hunger, safety, curiosity, etc) that constrain what will be learned from each experience. Nevertheless, animals develop a far better understanding of the world than modern reasoning models [97, 72], in part due to a remarkable intuition for how abstract variables interact.
To address this shortcoming, we must be able to differentiate meaningful intelligence from that which is trivial. [18] defines intelligence as ’skill-acquisition efficiency’, emphasizing the capacity to build upon prior knowledge. This relies on the ability to recognize or propose an unseen combination of elements as a novel solution, known as compositional generalization. In this paper, we formalize compositional generalization with abstract variables as the elemental units, using the principle of Interaction Information– mutual information among three or more random variables [108]. Interaction information has been used to understand sensory modality integration [74] and causal inference [30], but has not yet been put into a Reinforcement Learning (RL) framework. We bridge this gap by unifying Interaction Information with Semantic Information– the subset of information that is meaningful for survival [50]. To exemplify the relationship between Semantic Information and Interaction Information, consider an agent with two possible actions (i.e. causal interventions), where reward $r$ is determined by an exclusive or (XOR) operation between the agent’s action $a$ and a hidden (i.e. latent) binary variable $c$. XOR demonstrates Interaction Information in its purest form [80]: Action $a$ and reward $r$ are independent unless the agent knows the value of auxiliary variable $c$. We consider such variables as contextual since they mediate the dependencies between action, observations and reward, while not being directly causal. Here, we introduce Semantic Interaction Information (SII)– the contribution of Interaction Information between latent variables (and observables) to inference of the optimal action. More generally, SII emerges when conditioning on observables changes the mutual information between contextual variables and the optimal action. Conceptually, SII measures how much more certain an agent becomes about its goal by accounting for latent variable interactions, rather than treating them in isolation. Moreover, when variables interact in a systematic manner, learning can extrapolate to novel combinations of latent variables.
In summary, we argue that meaningful intelligence can be made formal, through Information Theory and RL, as learning from inference in environments with compositional latent structure (often studied under the umbrella of meta-learning [86, 38, 51]). Our support for SII as a mediator of meaningful intelligence can be split into three primary sections: (1) Cognitive Gridworld (CG): A minimal task that rewards agents for performing mental navigation in a compositional latent space. (2) Representation Classification Chains (RCCs): A novel architecture that consolidates experience into parametric embeddings to be used as building blocks for future learning. (3) Offline Optimization: A demonstration of the usefulness of RCCs for zero-shot control in novel spaces.
2 Related Works
2.1 Compositional Generalization
Learning to interact with the environment was originally described by Classical Behaviorism [92], and was formalized in RL as the reinforcement of actions that maximize internal reward (conditioned on the state of the world) [9, 94, 82]. Generalization in unambiguous environments can then be described as learning how independent dimensions of the state space interact in a manner that extrapolates to novel situations [34]. This capacity to recognize a novel combination of elements as a coherent whole is referred to as compositional generalization [47, 52]. Compositional representations emerge naturally in recurrent networks trained to perform complex sensory predictions [10]. Other works explore how such representations may be formed by specialized regions of the brain. The Tolman-Eichenbaum Machine (TEM) [110] models the Hippocampus as a network that learns to navigate a grid of past experiences. The ”internal action” of navigating between sensory observations is proposed to convey their relationships and facilitate prediction. Similarly, models of animal behavior and perception describe sequential inference akin to navigation towards a ”decision boundary” [91, 31, 48, 57, 12, 22, 45, 71, 84, 85]. This suggests that the internal ’commitments’ of a decision-making process can be treated as elemental units, allowing for compositional generalization to novel combinations of relevant decision-variables. A minimal task which allows researchers to design and manipulate hidden interactions could be a valuable tool for general intelligence research, complementary to the modern, large-scale benchmarks currently used [17, 78, 32]. The Cognitive Gridworld is an attempt to provide a formally grounded setting which explores the challenges of compositional generalization when the world is not fully observable.
2.2 Partially Observable Markov Decision Processes
Animals often learn without the need for concrete examples or symbols to reason with. For instance, riding a bike can become automatic without an understanding of all its mechanisms. Such abstract reasoning is often implicit—a ’gut-feeling’ that is effortlessly utilized but not as easily explained. Inference under such uncertainty can be modeled with a POMDP [63, 43]. POMDPs utilize a belief-state, a probability distribution over possible world states which is updated either passively by observation or actively via goal-directed exploration [39, 46]. Distributional representations suffer from the curse of dimensionality, so effective agents must encode goal-relevant information into a compressed approximation of the belief-state [5, 81]. Incidentally, belief-like representations have been implicated in the dopaminergic neurons involved in RL [36, 35]. Furthermore, cortical dynamics are often found to lie on a low dimensional manifold [56, 20]. In this work, we consider the compositional generalization that emerges from the dynamics of approximate inference. Observations in the Cognitive Gridworld are fundamentally ambiguous, without any inherent spatiotemporal structure that could circumvent the need to learn to approximate the underlying world model.
2.3 Generative modeling
The brain is hypothesized to cluster observations by independent latent causes [29, 25] through mechanisms such as attention and computation through dynamics [66, 105, 103]. Together, these form a traversable cognitive map [101, 8, 28]—an abstract space of task-relevant variables. This capability is believed to emerge via meta-learning [53, 107], wherein agents learn a repertoire of computations that can be reconfigured as needed for flexible inference [113, 102, 11]. Modern meta-learning algorithms utilize conditional generative models to produce data for the agent to train on. For instance, PEARL learns a posterior over contextual variables that summarize recent experience, and conditions the agent on samples from this posterior to facilitate few-shot acquisition of novel goals [77]. Additionally, [2] show that diffusion-based decision-making can generalize by combining conditioning variables at inference time. [1] consider the task of continual learning to be identifying generative factors that are shared between datasets. Using this assumption, their VASE algorithm learns independent factors in the data and recombines them to generate augmented data to train on. Such research illustrates the value of a simple, standardized meta-learning task centered around compositional abstractions. To demonstrate the usefulness of the Cognitive Gridworld, we show that the manner by which latent variables interact can be consolidated into embeddings. Learning these compositional embeddings, which encode how a variable will interact with others, requires conditioning a generative model on the inference model, and vice versa. To achieve this, we develop Representation Classification Chains and show that, after learning, the process can be inverted to evaluate locations on a novel cognitive map, with respect to preferred observations.
2.4 Setup
Our goal is to facilitate the study and development of agents that can generalize from goal-directed inference. To simplify our scope, we assume that for any specific goal, only a small number of latent variables are relevant, and that they interact in a systematic and learnable manner. We refer to the actual value of a latent variable as its realization, which is governed by a random process wherein each variable, in a context of $C$ relevant variables, takes on a value from $R$ possible realizations:
$\mathbf{r}\overset{\text{i.i.d.}}{\sim}U(\{1,\dots,R\})^{C},\quad\text{where }r_{c}\in\mathbf{r}\text{ is the realization of variable }c.$
The agent observes the world through a $d_{o}$-dimensional observation space $\mathcal{O}$ (Figure 1). At time $t$, the probability of each independent observable $o^{i}\in\{0,1\}$ within observation $\mathbf{o}_{t}$ depends on realizations $\mathbf{r}$ and is parameterized by variable interactions $Z$. In total, there are $|Z|=d_{o}\times C\times(C-1)$ pairwise and independent interactions. We let $\ell_{\mathbf{z}^{i}}(\mathbf{r})$ denote the likelihood $P_{\mathbf{z}^{i}}(o^{i}=1\mid\mathbf{r})$, which forms a tensor of order $C$ with $R$ entries along each axis. Each interaction $z^{i}_{cc^{\prime}}\in\mathbf{z}^{i}$ (the modulation from variable $c^{\prime}$ onto $c$) shifts the phase of the likelihood with respect to $r_{c}$, functionally ”rolling” values of $\ell_{\mathbf{z}^{i}}(\mathbf{r})$ along axis $c$ (see supplementary Figure 9). To generate Interaction Information in an interpretable manner, we model $\ell_{\mathbf{z}^{i}}(\mathbf{r})$ around the exclusive or operation via a procedure detailed in subsection A.1 (customize-able to any arbitrary function of $Z$). Importantly, these joint likelihoods cannot be factorized without losing information. Finally, an episode $E$ is defined by the tuple $(\mathbf{r},g,Z,\mathbf{o}_{1:T})$, comprising the realizations, a random goal index $g\sim\{1,\dots,C\}$, the interactions, and a trajectory of $T$ i.i.d. observations.
The agent’s objective is to infer the realization of goal variable $r_{g}$ and select optimal action $a^{\star}=r_{g}$ from $a\in\{1,\dots,R\}$. Optimal inference of $r_{g}$ can be achieved with sequential Bayesian updating:
$P_{Z}(\mathbf{r}\mid\mathbf{o}_{1:t})\propto P_{Z}(\mathbf{o}_{t}\mid\mathbf{r})\,P_{Z}(\mathbf{r}\mid\mathbf{o}_{1:t-1}),$
where the joint likelihood of all (independent) observables follows a geometric distribution:
$P_{Z}(\mathbf{o}_{t}\mid\mathbf{r})=\prod_{i=1}^{d_{o}}\ell_{\mathbf{z}^{i}}(\mathbf{r})^{o_{t}^{i}}(1-\ell_{\mathbf{z}^{i}}(\mathbf{r}))^{1-o_{t}^{i}}.$
Figure 1: Environment schematic for $C=2$. Observations $\mathbf{o}_{t}$ are generated stochastically. variable interactions $Z$ parameterize the likelihood $P_{Z}(\mathbf{o}\mid\mathbf{r})$ and variable realizations $(r_{1},r_{2})$ fix the probability of sampling each observable $o^{i}$.
Thus, while the real world involves a vast number of latent variables, the agent’s goal effectively reduces the world to a context of $C$ relevant variables—specifically, the goal variable itself and the contextual variables that generate SII. In other words, although the state space may be high dimensional, the optimal policy likely lies on a smaller $C$-dimensional manifold at any given time.
Our results are structured as follows: (i) Demonstrating how interactions impact Bayesian observers; Then training (ii) RNNs to capture Semantic Interaction Information; (iii) RCCs to learn how variables interact; (iv) RL-agents to identify regions of latent space that generate preferred observations.
3 Results
3.1 Semantic Interaction Information mediates the relative accuracy of Bayesian observers in a stationary POMDP
A natural baseline for our setting is a Bayesian observer whose marginal beliefs ($B_{tc}$) are updated independently. The likelihood for this Naive observer is acquired by marginalizing the joint likelihood (Figure 2a). Semantic Interaction Information quantifies the contribution of Interaction Information to marginal beliefs as the kullback-leibler divergence ($\mathcal{D}_{\mathrm{KL}}$) between $B_{tg}$ acquired through Joint and Naive Bayesian inference:
$\text{SII}_{t}=\mathcal{D}_{\mathrm{KL}}\!(B_{tg}^{\mathrm{Joint}}\|B_{tg}^{\mathrm{Naive}})$
(A formal derivation is provided in subsection A.2). Figure 2b compares the posterior distributions of the Bayesian observers. We observe that, for both Joint and Naive Bayes, the average performance (calculated as $\langle 1-[\arg\max_{r}\bigl(B_{tgr}\bigr)=a^{\star}]\rangle_{E}$ using Iverson bracket notation) improves with the number of evidences received and declines with the number of interactions. However, the relative accuracy attained by joint inference when compared to naive inference grows rapidly with the number of interactions, in accordance with the Semantic Interaction Information. In summary, the Cognitive Gridworld demonstrates that even in stationary POMDPs, systematic gaps emerge between nuanced and Naive Bayesian observers, despite both possessing perfect memory.
Figure 2: The cost of Naive Bayes grows with time and interactions. (a) Example Joint (matrices) and marginalized (vectors) likelihoods. (b) Top: Accuracy of Joint (left) and Naive (right) Bayes across varying context sizes. Bottom: Relative accuracy (left) and Semantic Interaction Information (right). Circles mark four equidistant reference time-points throughout inference.
3.2 Experiment 1: The Cognitive Gridworld requires flexible neural dynamics
The Cognitive Gridworld inverts the standard dependency structure of most tasks. Typically, independent features of an observation space (e.g., pixels or tokens) become conditionally dependent given the realization of a latent variable (e.g., a label or concept). In a CG, conditioning on latent variable realizations causes observables to become independent. The lack of spatial or temporal correlations in the observation space yields a uniform distribution of data, such that all learnable structure is latent. The network performing sequential inference, referred to as the Classifier, consists of Long-Short-Term-Memory (LSTM) units and receives $Z$ and $\mathbf{o}_{1:T}$. To demonstrate that arbitrary neural dynamics that encode $Z$ and $\mathbf{o}_{1:T}$ are insufficient, we benchmark a Fully Trained RNN against reservoir computing (Echo State RNN). While Echo State dynamics provide random temporal features on which to learn a function of recent inputs, they suffer from fundamental limitations that may be viewed as a tradeoff between discrimination and generalization, memory and nonlinearity or recall and precision [3, 21, 6, 40, 99, 98]. As a result, we predicted that the performance of an Echo State network would resemble that of a Naive Bayesian observer.
3.2.1 Interaction Information underlies network capability at inference time
To ensure the Echo State network has a rich teaching signal to learn from, we train with the optimal belief as the target (see subsection 5.1 for details), and to ensure performance does not reflect memory limitations, the network outputs belief updates $\Delta B_{t}\in\mathbb{R}^{C\times R}$ which we accumulate over time (Figure 3a). In Figure 3b we observe that the performance of the Fully Trained network is explained by Joint Bayes, meanwhile, as predicted, the Echo State network reaches exactly the accuracy of Naive Bayes (Figure 3c). Crucially, the Echo State network performs optimally in the case of a single latent variable ($C=1$), where joint inference is equivalent to independent inference. This indicates the sub-optimality at $C=2$ is not due to basic memory limitations, but a specific inability to approximate joint inference. Moreover, the relative accuracy between the Fully Trained and Echo State networks mirrors the exact ratio seen between Joint and Naive Bayes (Figure 3d). These results suggest that task-aligned dynamics [90] are particularly necessary when latent variables interact. A comparison with $C=3$ can be seen in supplementary Figure 16a.
Figure 3: Recurrent Neural Networks align with theoretical predictions. (a) Architecture (left) and gradient flow (right) of the Classifier. Only the goal belief-state receives a gradient. (b-d) Same as Figure 2b (for $C=1,2$) with Fully Trained and Echo State Networks. Markers indicate four equidistant reference time-points throughout inference.
3.2.2 Naive Bayes predicts the failure mode exhibited by Reservoir Computing
A closer look at the belief dynamics in $C=2$ in Figure 4 reveals an unexpected failure mode analogous to hallucinations: high confidence in an incorrect conclusion. Figure 4a shows a representative trajectory of goal belief under Joint and Naive Bayes (additional trajectories are provided in supplementary Figure 10). In Figure 4b, we pool the belief assigned to the correct action (i.e. the probability of a hit) over episodes and analyze its density at each step of the trajectory. Early in the trajectory (blue), the distribution of hits is centered near chance. As time progresses, the density shifts rightward, reflecting a growing probability of a hit. Unexpectedly, for models that ignore SII (the Echo State network and Naive Bayes), a second, disjoint peak emerges below chance (Figure 4b, right column). This peak represents a significant fraction of trajectories where misinterpretation of evidence lead the agent to become wrong and confident. In other words, the consequence of neglecting SII is not just a reduction in accuracy, but the active creation of confident, incorrect belief. Therefore, in the presence of Interaction Information, model confidence becomes unreliable.
Figure 4: Failure to capture SII can induce hallucinations. (a) Sequential updating of example posteriors under Joint and Naive inference. (b) Distribution of hits and misses at each step, pooled over episodes. Misinterpreting evidence yields episodes with performance below chance.
3.3 Experiment 2: Learning interactions requires variational inference
Thus far, we have established how Interaction Information can impede optimal action selection. While networks in Experiment 1 were trained with explicit access to the true interactions $Z$, subsequent experiments assume these latent interactions are unknown but learnable. To achieve this, we introduce compositional embeddings that encode how variables interact. Consider a large set $\mathcal{S}$ of latent variables, where each individual variable (indexed by $s\in\mathcal{S}$) represents a subset of mutually exclusive realizations of the world. These may be the seasons or the time of day, for example. For a specific goal, like waking up on time, and given the observed sunlight, most other latent variables are irrelevant. However, if the interaction between season and time correlates with how the season would interact with another latent variable, possibly relevant to a future goal, this relationship is encoded in season’s compositional embeddings. Formally, we associate variable $s$ with a set of learnable embeddings $\{(\mathbf{k}_{s}^{i},\mathbf{q}_{s}^{i})\}_{i=1}^{d_{o}}\in\mathcal{E}$. The context, i.e. subset of relevant variables $C\subset\mathcal{S}$, has interactions $z_{cc^{\prime}}^{i}=(\mathbf{q}_{c}^{i})^{\top}\mathbf{k}_{c^{\prime}}^{i}$ for all distinct pairs $c,c^{\prime}\in C$ (Figure 5).
Figure 5: Compositional embeddings are learned indirectly via goals. Schematic demonstration of compositional generalization in latent space. Training episodes (top) contain at most one testing variable, which is never the goal (green). Testing episodes (bottom) consist entirely of testing variables. Success requires testing variable embeddings to be learned through their implicit relationships to training goals. This setup evaluates zero-shot generalization to novel contexts and novel goals.
[…]
3.3.1 Naturalistic feedback is sufficient for learning interactions
Inspired by the V-JEPA 2-AC architecture [4], we separate the problem into embedding and prediction. Modules of our Representation Classification Chains (RCCs) (Figure 6b) bootstrap off each other without sharing gradients: a Classifier uses $\hat{Z}$ to predict realizations $\hat{\mathbf{r}}$, while a Generator uses $\hat{\mathbf{r}}$ to predict observations (see subsection 5.2). To train the Classifier we treat the goal belief as a policy, sampling a single action and using a cross-entropy-style loss to provide feedback that is sparse (sampled at the end of the trajectory), partial (goal-variable only), and binary (reward/punishment). Concurrently, the Generator network is trained via a self-supervised loss and the gradient flows from $\hat{Z}$ to $\hat{\mathcal{E}}$, teaching the embedding space to be useful for prediction. Figure 6c illustrates testing episode accuracy throughout training, demonstrating that $\hat{\mathcal{E}}$ is indeed learnable from naturalistic feedback.
Figure 6: A variational architecture learns compositional embeddings from reward. (a) Relevant variables interact via learnable embeddings to form interactions. (b) Forward pass and gradient flow of the Classifier and Generator. The Classifier learns from rewards while the Generator uses self-supervised-learning (SSL). (c) Testing episode accuracy of the Classifier throughout training.
3.4 Experiment 3: Latent space navigation enables offline optimization
[…]
$\mathcal{R}(\mathbf{r},\mathbf{\Omega},\hat{\ell},\hat{Z})=\exp\left(\frac{1}{d_{o}}\sum_{i}\bigl(\ln(\hat{\ell}_{\hat{\mathbf{z}}^{i}}(\mathbf{r}))\Omega^{i}+\ln(1-\hat{\ell}_{\hat{\mathbf{z}}^{i}}(\mathbf{r}))(1-\Omega^{i})\bigr)\right).$
[…]
Figure 7: Conditional generative modeling enables optimization in compositional spaces. (a) Schematic illustrating the mapping of preferred observations ($\mathbf{\Omega}$) to their respective likelihoods and the cumulative landscape (accumulated over $i$ in subsection 3.4). (b) An example traversal, from the lowest to the highest point on the landscape, changes observations to best match the agent’s preference. (c) Controller learning curves (mean $\pm$ SEM, $n=20$ initializations of $\mathbf{\Omega}$). Performance is evaluated as $\mathcal{R}(\text{argmax}_{\mathbf{r}}\pi(\mathbf{r}),\mathbf{\Omega},\ell,Z)/\max_{\mathbf{r}}\mathcal{R}(\mathbf{r},\mathbf{\Omega},\ell,Z)$ using the ground truth $\ell$ and $Z$.
The RL agents are trained through a standard Advantage Actor-Critic (A2C) algorithm. We compare Controller training across three conditions that vary in how they estimate $\hat{Z}$ and $\hat{\ell}$:
[…]
Training fully offline with the generative model achieves the similar performance and efficiency as training offline with direct access to the joint likelihood and training online with direct interactions with the environment (Figure 7c). It is worth noting that the rapid learning seen in Figure 7c required a large batch of 20,000 trajectories per episode, each followed by an interaction with the environment in the Online condition. In the Offline conditions, training did not require any interactions with the environment. Figure 8 displays several examples of the Offline w/ Generator agent’s deterministic policy ($\text{argmax}_{\mathbf{r}}\pi(\mathbf{r})$) throughout several learning trajectories. These traversals reveal the utility of compositional embeddings for offline optimization. More representative landscapes and trajectories can be found in supplementary Figure 15.
Figure 8: Example offline learning trajectories w/ Generator. The evolution of the deterministic policy, $\text{argmax}_{\mathbf{r}}\pi(\mathbf{r})$, is plotted throughout offline training from initialization (red circles) to the end of training (green stars). Trajectories are overlaid on the preference landscapes to demonstrate navigation through an internal Cognitive Gridworld.
4 Discussion
In this work, we attempted to formalize the ability to generalize from past experience as an instance of variational inference with latent variables that interact in a parametric embedding space. We introduced the concept of Semantic Interaction Information to quantify the reward-relevant information that emerges from these latent variable interactions. To operationalize interactions, we developed the Cognitive Gridworld, a modular environment designed to differentiate meaningful and trivial intelligence. Crucially, Semantic Interaction Information not only explained task performance on average, but also predicted a failure mode reminiscent of hallucinations. These findings support the treatment of Naive Bayes as a theoretical lower bound on meaningful intelligence in stationary POMDPs. While the exponential growth of potential interactions renders exhaustive learning computationally intractable, goals create the necessary pressure to discover specific relationships.
Building off previous work, we demonstrated that Representation Classification Chains enable an agent to learn variable embeddings from natural feedback and passive experience. This novel architecture considers continual meta-learning as the conjunction of variational and sequential inference, continually learning embeddings on a separate timescale from the dynamics of inference. The learned embeddings can then be utilized to optimize novel objectives without further environmental interaction. It is notable that while Representation Classification Chains were trained on individual training variables, compositional embeddings enabled the Offline Controller to optimize a policy in a multi-dimensional joint action space of testing variables. This extends the work of [2], which cast model-free RL as a conditional generative modeling problem to enable compositional generalization in the input space. Conversely, by recasting latent space traversals as a model-free RL problem, RCCs generalize to novel combinations of outputs (e.g. decision variables, control variables etc). This may be particularly helpful for optimizing non-differentiable objectives. Future work should attempt to integrate RCCs with the VASE algorithm to identify which embeddings are relevant to a task, rather than assuming this is known.
Beyond artificial intelligence, we also aspired to provide a compelling model of meaningful intelligence in animals. Specifically, Semantic Interaction Information offers a unifying lens through which to interpret the diverse empirical findings on the Orbitofrontal Cortex (OFC), a region critical for cognitive flexibility and highly specialized in primates [15, 23, 88, 58, 19, 37, 41, 83, 73]. This area is necessary for latent state inference, adaptation to context, and evaluation of feature conjunctions [24, 95, 26, 89, 16, 70, 93, 61, 87]. Furthermore, OFC activity is thought to represent a multidimensional abstract space with reward-dependent compression, encoding goals and a “common currency” value, including the value of information itself and of imagined outcomes [68, 69, 44, 95, 7, 64, 75, 14]. These seemingly disparate functions—inference, evaluation, imagination, and compositionality—all follow naturally from the objective of maximizing SII.
This view also offers a framework for understanding the complementary cognitive maps of the Hippocampus and OFC [112, 111, 49, 106, 62, 54]. While the Hippocampus is well suited for dynamic environments where order matters [76], sequential Bayesian inference in a stationary environment is permutation invariant. According to our framework, the Hippocampus compresses ordered sensory sequences into embeddings which encode how underlying causal factors would interact if they co-occurred. Once consolidated into the cortex, the task-relevant embeddings must be integrated across multiple timescales, sensory modalities and levels of abstraction. This likely requires top-down control of inter-cortical communication. We propose that the OFC is this cortical conductor, recruiting the distributed cortical representations of independent latent variables into a brain-wide model of the world through its robust connections with limbic, thalamic, and neuromodulatory systems [79, 27, 96, 55, 42, 65, 109]. This would imply that the integration of information distributed throughout the cortex should be permutation invariant. Incidentally, treating the environment as fixed (yet unknown) is precisely the assumption that makes meta-learning tractable [67]. In summary, we propose that maximization of Semantic Interaction Information serves as an objective for which the OFC and Hippocampus collaborate; the Hippocampus consolidates ordered sequences into compositional representations embedded throughout the cortex, meanwhile the OFC expands interactions between relevant representations into a cognitive map for optimizing novel objectives.
5 Methods
5.1 Experiment 1:
Network goal belief.
The goal belief $B^{\text{Net}}_{tgr}$ is extracted from the accumulated output $\mathcal{M}_{tcr}=\sum_{t^{\prime}=1}^{t}M_{t^{\prime}cr}$ (where $M_{t^{\prime}cr}$ is single timestep network pre-activations) via a softmax over possible realizations $B^{\text{Net}}_{tgr}=\frac{e^{\mathcal{M}_{tgr}}}{\sum_{r^{\prime}}e^{\mathcal{M}_{tgr^{\prime}}}}.$ Although observations are conditionally independent given both latent variable realizations $(r_{g},r_{c})$, the optimal maximum a posteriori (MAP) estimate of $r_{g}$ requires marginalizing the joint likelihood over $r_{c}$ before maximizing:
$\arg\max\limits_{r_{g}}P_{Z}(r_{g}\mid\mathbf{o}_{1:T})=\arg\max\limits_{r_{g}}\ln\left(\sum_{r_{c}}\exp\left(\sum_{t=1}^{T}\ln P_{Z}(\mathbf{o}_{t}\mid r_{g},r_{c})\right)\right).$
Crucially, because the summation over $r_{c}$ envelops the temporal accumulation, the log-likelihood cannot be decomposed into a simple iterative summation over time for the marginals. Therefore, sequential Bayesian updating is only Markovian with respect to an intractable joint posterior $P_{Z}(r_{g},r_{c}\mid\mathbf{o}_{1:t})$. Since $\Delta B_{tc}$ is restricted to marginal beliefs to remain tractable, re-evaluating the marginal $B_{tc}$ in light of a new observation $\mathbf{o}_{t+1}$ will not be Markovian. Therefore, performing this update accurately requires the network’s memory state to encode the missing joint dependencies—specifically, the interaction information—that is lost when only tracking marginals.
Classifier loss.
To provide a rich teaching signal, we minimize a symmetrized KL divergence between the network and optimal goal belief-state (where $\alpha=0.5$ gave the best results):
$\mathcal{L}_{\text{SII}}=\left(\frac{1}{2}\mathcal{D}_{\mathrm{KL}}(B_{tg}^{\mathrm{Joint}}\,\|\,B_{tg}^{\mathrm{Net}})+\frac{1}{2}\mathcal{D}_{\mathrm{KL}}(B_{tg}^{\mathrm{Net}}\,\|\,B_{tg}^{\mathrm{Joint}})\right)^{\alpha}.$
5.2 Experiment 2:
Classifier loss.
At the end of each trajectory we sample an action $a\sim B_{Tgr}$ and receive outcome $\delta(a-r_{g})\in\{0,1\}$. We train using a cross-entropy-style objective with an entropy bonus ($\beta_{1}$):
$\mathcal{L}_{\text{Classifier}}=\dfrac{-1}{T}\sum_{t}\underbrace{\delta(a-r_{g})}_{\text{rewarded}}\ln B_{tga}+\underbrace{\bigl(1-\delta(a-r_{g})\bigr)}_{\text{unrewarded}}\ln(1-B_{tga})-(-\beta_{1}\underbrace{B_{tga}\ln B_{tga}}_{\text{entropy bonus}}),$
Gradients from $\mathcal{L}_{\text{Classifier}}$ update only the Classifier parameters.
Generator loss.
We sample a latent configuration $\hat{\mathbf{r}}$ from the Classifier belief at $t=T$ (without a gradient), and train the Generator to maximize the likelihood of observations. Since the true generative distribution is unknown, we construct an empirical target distribution $\tilde{P}(\mathbf{o})$, where the likelihood of each independent observable is given by its trajectory average $\tilde{p}_{Z}(o^{i}=1)=\langle o^{i}\rangle_{t}$. We use this to compute a symmetric $\mathcal{D}_{\mathrm{KL}}$ (hence a self-supervised loss):
$\mathcal{L}_{\text{Generator}}=\frac{1}{2}\mathcal{D}_{\mathrm{KL}}\bigl(\tilde{P}(\mathbf{o})\,\|\,\hat{P}_{\hat{Z}}(\mathbf{o}\mid\hat{\mathbf{r}})\bigr)+\frac{1}{2}\mathcal{D}_{\mathrm{KL}}\bigl(\hat{P}_{\hat{Z}}(\mathbf{o}\mid\hat{\mathbf{r}})\,\|\,\tilde{P}(\mathbf{o})\bigr)+\mathcal{L}_{\text{reg}}.$
Gradients from $\mathcal{L}_{\text{Generator}}$ update the embeddings via gradient flow through $\hat{Z}$.
Gradient stabilization (soft clip).
To avoid vanishing gradients when clipping probabilities, we use a differentiable soft clip:
$\text{soft clip}(x,\epsilon_{clip})=\epsilon_{clip}+\sigma\left(\frac{\hat{x}-\epsilon_{clip}}{\epsilon_{clip}}\right)(\hat{x}-\epsilon_{clip}),\quad\hat{x}=(1-\epsilon_{clip})-\sigma\left(\frac{1-\epsilon_{clip}-x}{\epsilon_{clip}}\right)(1-\epsilon_{clip}-x).$
We found that training reliability to improve dramatically when $\epsilon_{clip}=1e-3$.
5.3 Experiment 3:
Controller loss.
The Controller is trained using an Advantage Actor-Critic (A2C) algorithm to maximize the intrinsic reward. The Actor and Critic both receive $\hat{Z}$, and output a policy over $\mathbf{r}$ and the expected $\mathcal{R}$, respectively. At each episode, $\mathbf{r}$ is sampled stochastically from the Actor’s policy $\pi(\mathbf{r})$. In contrast to the Classifier, whose outputs are marginal beliefs, the Actor’s policy is a probability distribution over the full action space so $\mathbf{r}$ is sampled from all dimensions jointly. The Critic predicts the intrinsic reward V(Z^)V($\hat${Z}). The advantage $A$ is calculated as the difference between the true intrinsic reward and the Critic’s prediction:
[…]
The Controller loss $\mathcal{L}_{cntrl}$ integrates the Critic’s prediction error, the Actor’s policy gradient and a policy entropy bonus weighted by $\beta_{cntrl}$ to encourage exploration:
$\mathcal{L}_{cntrl}=A^{2}-\ln(\pi(\hat{\mathbf{r}}\mid\hat{Z}))A-\beta_{cntrl}(-\pi\ln\pi)$
where the entropy bonus scaling $\beta_{cntrl}$ decays over the training episodes.
Acknowledgments
I would like to express my gratitude to Eran Lottem, Jonathan Kadmon, Jan Bauer and Vladimir Shaidurov, for constructive discussion and support.
5.3.1 Data availability
The full code is available on the Cognitive Gridworld Github.
Appendix A SUPPLEMENTARY
A.1 Generative process details
Figure 9: A flexible process for embedding Gridworld structure into latent space. The embeddings of relevant variables are compressed into interactions which are then expanded to a discrete probability distribution over possible realizations of the world. The full process consists of first (i) compressing embedding vectors to their scalar interactions. Then (ii) expanding pairwise interactions to pairs of vectors, expanding the vector pairs to matrices through an outer product, and (for $C\geq 3$) expanding the matrices into a tensor by broadcasting each matrix over all other dimensions. Finally, a valid (iii) likelihood is ensured by a sigmoid transformation for normalization.
Compositional embeddings are acquired by sampling $\mathbf{k}^{i}_{s}\sim\mathcal{N}(0,1)^{d_{\mathcal{E}}}\quad\mathbf{q}^{i}_{s}\sim\mathcal{N}(0,1)^{d_{\mathcal{E}}},\quad\forall s\in\mathcal{S}$, orthogonalizing across $\mathcal{O}$ via a Gram–Schmidt process and finally normalizing to 1 along $d_{\mathcal{E}}$. In each Episode E, the generative process is described by:
$o^{i}\sim P(o^{i}=1\mid\mathbf{k}^{i}_{1},\mathbf{q}_{{1}}^{i},\dots,\mathbf{k}^{i}_{{C}},\mathbf{q}_{{C}}^{i},r_{1},\dots,r_{C})),\quad\forall i\in\mathcal{O},$
which can be shortened to:
$\quad\ell_{\mathbf{z}^{i}}(\mathbf{r})=\mathrm{generation}(\underbrace{\mathrm{compression}(\mathbf{k}^{i}_{{1}},\mathbf{q}_{{1}}^{i},\dots,\mathbf{k}^{i}_{{C}},\mathbf{q}_{{C}}^{i})}_{\underbrace{\mathrm{expansion}(z_{1}^{i},\dots,z^{i}_{C\times(C-1)})}_{\ell_{\mathbf{z}^{i}}}},r_{1},\dots,r_{C})$
Extrapolation to an out of distribution set $\{\mathbf{k}^{i}_{c},\mathbf{q}_{c}^{i},\dots,\mathbf{k}^{i}_{{C}},\mathbf{q}_{{C}}^{i}\}$ may occur if the resultant $\mathbf{z}^{i}$ falls within the manifold of training $\mathbf{z}^{i}$. The low rank structure of $\mathbf{z}^{i}\in\mathbb{R}^{C\times(C-1)}$ is embedded in joint likelihood, $\ell_{\mathbf{z}^{i}}\in[0,1]^{{R}^{C}}$. Below we step through the default expansion from $\mathbf{z}^{i}\to\ell_{\mathbf{z}^{i}}$ in 2 steps.
Step 1. Expanding $z_{cc^{\prime}}^{i}\to\mathbf{v}^{i}_{cc^{\prime}}$:
First, we expand each scalar $z^{i}_{cc^{\prime}}$ to vector $\mathbf{v}^{i}_{cc^{\prime}}$ through $\Theta$ and $\omega$:
$\mathbf{v}_{cc^{\prime}}^{i}(r)=\sum_{n=0}^{N-1}\Theta(n,r)\ \omega(n),\quad N=1+2R.$
Here, $\Theta$ is a cyclic shift of sinusoidal base values:
$\Theta(n,r)=\lambda\sin\!\biggl(\frac{2\pi}{N}\bigl(((n-r)\bmod N)-N\bigr)\biggr),\quad\lambda\ \text{= likelihood temperature},$
and $\omega$ gives softmax weights over circular distances from $z_{cc^{\prime}}^{i}$:
$\omega(n)=\mathrm{softmax}_{n}\!\left(-\left[\frac{N}{2\pi}\cdot\min\!\big(|\theta(n)|,\ 2\pi-|\theta(n)|\big)\right]^{2}\right),\quad\text{where}\quad\theta(n)=2\pi\left(\frac{n}{N}-z_{cc^{\prime}}^{i}\right).$
Functionally, $z_{cc^{\prime}}^{i}$ performs a smooth approximation of the discrete roll operation of a sinusoid. Note that $N>R$, $\mathbf{v}_{cc^{\prime}}^{i}$ is shorter than a full period, allowing for asymmetry in $\mathbf{v}_{cc^{\prime}}^{i}$.
Step 2. Expanding $\mathbf{v}_{cc^{\prime}}^{i}(\mathbf{v}_{c^{\prime}c}^{i})^{\top}\to\ell_{\mathbf{z}^{i}}\in[0,1]^{R^{C}}$:
Next, we expand outer products $\mathbf{v}_{cc^{\prime}}^{i}(\mathbf{v}_{c^{\prime}c}^{i})^{\top}$ to tensor $\ell_{\mathbf{z}^{i}}\in[0,1]^{R^{C}}$ through:
$\ell_{\mathbf{z}^{i}}\;=\;\begin{cases}\sigma\!\big(\mathbf{v}_{cc}^{i}\big),&C=1,\\[6.0pt] \sigma\!\left(\sum_{\begin{subarray}{c}\{cc^{\prime}\}\\ c\neq c^{\prime}\end{subarray}}E_{cc^{\prime}}\!\Big(\mathbf{v}_{cc^{\prime}}^{i}(\mathbf{v}_{c^{\prime}c}^{i})^{\top}\Big)\right),&C\geq 2,\end{cases}$
where $\sigma$ denotes element-wise sigmoid and $E_{cc^{\prime}}:\mathbb{R}^{R\times R}\to\mathbb{R}^{R^{C}}$ is an embedding operator that expands $\mathbb{R}^{R\times R}$ to $\mathbb{R}^{R^{C}}$ by adding singleton dimensions along all axes but $c,c^{\prime}$ of $\ell_{\mathbf{z}^{i}}$. For example, for the case of $C=3:$
$\ell_{\mathbf{z}^{i}}=\sigma\bigl(\{\mathbf{v}^{i}_{12}(\mathbf{v}^{i}_{21})^{\top}\}_{12}+\{\mathbf{v}^{i}_{13}(\mathbf{v}^{i}_{31})^{\top}\}_{13}+\{\mathbf{v}^{i}_{23}(\mathbf{v}^{i}_{32})^{\top}\}_{23}\bigr).$
Important properties of likelihood constructions.
We chose our manner of expansion to ensure several properties:
  • Low rank structure: Broadcasting $\mathbf{v}_{cc^{\prime}}^{i}(\mathbf{v}_{c^{\prime}c}^{i})^{\top}\in\mathbb{R}^{R\times R}$ to $\mathbb{R}^{{R}^{C}}$ constructs the likelihood entirely from pairwise interactions, embedding low-rank structure into the tensor.
  • Constrained scale: Using a sine function for $\mathbf{v}_{cc^{\prime}}^{i}$ creates a linear (as a function of $C$) bound on the standard deviation of (pre-activation) elements of $\ell$.
  • Visual Interpretability: Constructing $\mathbf{v}^{i}_{cc^{\prime}}$ with a smooth function ensures local similarity in $P_{Z}(\mathbf{o}\mid\mathbf{r})$ such that similar vectors $\mathbf{r}$ generate similar observations.
  • Consistency: Using $z_{cc^{\prime}}^{i}$ to modulating phase, rather than scale, ensures different contexts and dimensions of $\mathcal{O}$ generate similar amounts of information.
  • Useful Baseline: Ensuring the potential for asymmetry in $\mathbf{v}^{i}_{cc^{\prime}}$ ensures that naive inference can perform above chance.
Future Directions of the Generative Process.
A natural extrapolation would be to construct $\mathbf{v}^{i}_{cc^{\prime}}$ using Fourier series, perhaps adding another dimension to $K$ and $Q$ for frequency specific interaction. Such a design would embed structure at multiple spatial frequencies while staying interpretable, similar to modules of grid cells in the Hippocampus [13]. Another natural extension would be to introduce sparsity to $K$ and $Q$, to move away from the current regime where all dimensions of $\mathcal{E}$ have some degree of interaction. In summary, 2 promising future directions:
  • Construct $\ell$ through inverse DFT (rather than a single frequency) for greater expressivity.
  • Sparsify embeddings by masking elements in $\mathcal{E}$ for greater interaction specificity.
A.2 SII Relation to Interaction Information
Interaction Information, $I(r_{g};r_{c};o^{i})$, quantifies how the mutual information between two random variables, $r_{g}$ and $r_{c}$, changes when conditioned on a third, $o^{i}$. Assuming the latent variables $r_{g}$ and $r_{c}$ are a priori independent ($I(r_{g};r_{c})=0$) (as they are in the setting we consider), the definition simplifies to (the negative of) conditional mutual information, or equivalently, the (negative) expected $\mathcal{D}_{\mathrm{KL}}$ between the conditional joint distribution and the product of its conditional marginals:
$I(r_{g};r_{c};o^{i})=-I(r_{g};r_{c}\mid o^{i})=-\mathbb{E}[\mathcal{D}_{\mathrm{KL}}(p_{Z^{i}}(r_{g},r_{c}\mid o^{i})\,||\,p_{Z^{i}}(r_{g}\mid o^{i})\otimes p_{Z^{i}}(r_{c}\mid o^{i}))]$
Note that the marginals of $p_{Z^{i}}(r_{g},r_{c}\mid o^{i})$ are not the posteriors of independent inference.
Let us now consider two types of Bayesian observers processing a sequence of observations $o^{i}_{1:T}$:
  1. 1.
    Optimal Observer: Updates its joint belief $B_{t}(r_{g},r_{c})$ correctly using $p_{Z^{i}}(o^{i}|r_{g},r_{c})$.
  2. 2.
    Naive Observer: Updates its joint belief $\tilde{B}_{t}(r_{g},r_{c})$ using a factorized likelihood
    $\tilde{p}_{Z^{i}}(o^{i}|r_{g},r_{c})\propto p_{Z^{i}}(o^{i}\mid r_{g})\otimes p_{Z^{i}}(o^{i}\mid r_{c})$
    because it is unaware of the conditional dependencies between $r_{g}$ & $r_{c}$.
For a single observation ($o_{1}$), the marginalized posteriors of both observers are identical:
$\sum_{r_{c}}p_{Z^{i}}(o^{i}_{1}|r_{g},r_{c})p(r_{g},r_{c})=\sum_{r_{c}}p_{Z^{i}}(o^{i}_{1}\mid r_{g})\otimes p_{Z^{i}}(o^{i}_{1}\mid r_{c})p(r_{g},r_{c})$
Therefore, for the Interaction Information produced by a single observation ($o^{i}_{1}$), it is valid to replace the conditional distributions with our Optimal and Naive observers:
$I(r_{g};r_{c};o^{i}_{1})=-\mathbb{E}[\mathcal{D}_{\mathrm{KL}}(B_{0}(r_{g},r_{c}|o^{i}_{1})\,||\,\tilde{B}_{0}(r_{g}|o^{i}_{1})\otimes\tilde{B}_{0}(r_{c}|o^{i}_{1}))]$
Thus, for one observation, Interaction Information is exactly the divergence between Optimal and Naive posteriors. However, for multiple observations, the inability of the Naive observer to encode Interaction Information in its belief can cause subsequent marginalized posteriors to diverge. Semantic Interaction Information (SII) quantifies the cumulative effect of this divergence on a task-relevant variable $r_{g}$. It’s defined as the expected $\mathcal{D}_{\mathrm{KL}}$ between the Optimal and Naive observers’ marginal belief over $r_{g}$ after $T$ observations:
$SII_{T}=\mathbb{E}[\mathcal{D}_{\mathrm{KL}}(B_{T}(r_{g}|o^{i}_{1:T})\,||\,\tilde{B}_{T}(r_{g}|o^{i}_{1:T}))]$
Equivalently, SII measures the residual Interaction Information in $\mathbf{o}_{1:T}$ after conditioning on $r_{g}$:
$SII_{T}\;\propto\;\mathbb{E}\!\Big[\mathcal{D}_{\mathrm{KL}}\!\big(p_{Z}(\mathbf{o}_{1:T}|r_{g})\;\|\;\prod_{t}\prod_{i}p_{Z^{i}}(o^{i}_{t}|r_{g})\big)\Big],$
revealing it as the cost of assuming that $r_{g}$ alone suffices to make observations i.i.d.
Summary: $SII$ is not $I(r_{g};r_{c};\mathbf{o}_{t})$, rather it is $I(\mathbf{o}_{1},\dots,\mathbf{o}_{T},r_{g})$, a measure of the accumulated, reward-relevant information which would be lost by neglecting interaction information. SII is the cost of a myopic focus on the goal-pertinent variable and distinguishes a genuine world model from strategies that simply exploit spurious correlations.
Figure 10: Additional examples of Bayesian inference for $C=2$. Representative examples of belief-state updating under Joint and Naive inference.
A.3 Learning Dynamics
Future theoretical work is still needed to specify the relationship between learning, dynamics and computation. For instance, we observed that while average performance improves throughout training for both Fully Trained and Echo State networks (Figure 11a-b), a negative correlation between accuracy and Semantic Interaction Information emerges early on (Figure 11c). It is important to highlight this critical finding that remains poorly understood.
Since the network initially knows nothing, the correlation between SII and accuracy begins at zero. Yet rapidly, failure becomes predictable from SII because the agent is failing in trickier trajectories that require second guessing Naive Bayes. Naivety may be conceptualized as a short-term reward which must be actively resisted for nuance, akin to an informational Marshmallow test [60]. Importantly, this correlation dissolves for the Fully Trained network, but persists when the recurrent weights are frozen. Despite the Echo State performance improving on average, its solution is moving further from optimality. A natural analogy is the exploration-exploitation tradeoff, where finetuning an early strategy could lead the agent away from discovering a superior strategy.
We hypothesize that this suboptimal strategy is stimulus bound [104]. To test this hypothesis, we investigate the Participation Ratio (i.e. dimensionality) of the Read-in activity and recurrent activity (Figure 11d). While a Fully Trained network expands the dimensionality of its recurrent dynamics, the Echo State network develops a more expressive encoder. Although this strategy improves accuracy, it may prevent the discovery of subtle changes in a trajectory that correspond to large jumps in latent space. Ultimately, we find that externally-driven dynamics are associated with a negative correlation between SII and accuracy. Despite early learning, the Echo State network remains likely to fail in trajectories with high SII, leading to an upper bound characteristic of Naive Bayes.
Figure 11: Early correlation between accuracy and SII predicts eventual performance ($C=2$). (a) Throughout learning, the testing accuracy at the final step of inference saturates to the performance of either Joint or Naive Bayes. (b) Final step testing accuracy during early training. (c) Correlation between Semantic Interaction Information (SII) and accuracy at the final step of inference. A negative correlation emerges in the beginning of training and persists for the Echo State network. (d) Difference in Participation Ratio (PR) between the recurrent and read-in activity.
A.4 Dis-entanglement
Misalignment between the Echo State network’s internal space and the true latent space could be explained mechanistically as entanglement. Entanglement in an RNN’s representation may emerge from the mixing of information by temporal correlations between neurons [59]. Therefore, we hypothesized that training only a read-in and read-out on an evidence accumulation task, which demands temporal correlations, would not only fail to overcome this entanglement, but would exacerbate it. Indeed, Principal Component Analysis (PCA) on the marginal beliefs for network and Bayesian agents was consistent with this prediction. When performing PCA on marginal beliefs after a single observation, all representations appeared factorized (Figure 12). Yet PCA on the marginal beliefs at the end of inference appeared entangled for the Echo State and Naive Bayes (Figure 13) which lose the ability to differentiate $r_{c}$ from $r_{c^{\prime}}$. Additionally, the Joint and Fully Trained representations allocate low entropy beliefs (i.e. high confidence) to all $(r_{c},r_{c^{\prime}})$ pairs. In contrast, the Echo State network and Naive Bayes represent confidence in an orthogonal direction to $r_{c}$ and $r_{c^{\prime}}$.
Figure 12: Belief representations are initially factorized. Top 2 Principal Components of the marginal beliefs after a single observation. Beliefs are colored by the realization sum ($r_{c}+r_{c^{\prime}}$), difference ($r_{c}-r_{c^{\prime}}$), and belief entropy ($\sum_{c}\sum_{r}B_{tcr}\ln B_{tcr}$). $R^{2}$ indicates the variance explained of the respective variables by the top 2 components.
Figure 13: Failing to capture interaction information causes entanglement over time. Same as Figure 12, for the final observation. The Fully Trained network maintains factorized representations akin to Joint Bayes. The Echo State and Naive representations of $r_{c}$ and $r_{c^{\prime}}$ become entangled.
For a mechanistic account of this entanglement, let $\mathbf{e}_{r}\in\{0,1\}^{R}$ represent a one-hot indicator vector which inquires about the probability of realization $r$. In $P_{z^{i}}(o^{i}\mid r)$, the mapping from scalar $z^{i}$ to a vector of $R$ updates is mediated by expansion $f(z^{i})$. For $C=1$, inquiring about a specific realization $r=r^{\prime}$ projects the indicator vector onto the likelihood via:
$P_{z^{i}}(o^{i}=1\mid r=r^{\prime})=\mathbf{e}_{r^{\prime}}^{T}f(z^{i}).$
Crucially, because different realizations of the same variable are mutually exclusive, the likelihood of a specific $r=r^{\prime}$ is perfectly isolated from all other potential realizations and $P_{Z}(r=r^{\prime}\mid\mathbf{o})$ may be updated independently. Next, consider a context of two latent variables ($C=2$). Here the interactions $Z^{i}=(z^{i}_{gc},z^{i}_{cg})$ map to a joint space via the bilinear tensor product $f(z^{i}_{gc})\otimes f(z^{i}_{cg})\in\mathbb{R}^{R\times R}$, forming an $R\times R$ matrix of belief updates. The bilinear construction generates synergy. In other words, the combination of variables loses predictive power if the variables are considered in isolation, i.e. the whole is greater than the sum of its parts [74, 6]. We can see this because evaluating the likelihood of $o^{i}=1$ given a realization $r_{g}=r^{\prime}$ now involves projecting indicator $\mathbf{e}_{r_{c}}$ onto a basis shared with $\mathbf{e}_{r_{g}}$:
$P_{Z^{i}}(o^{i}=1\mid r_{g}=r^{\prime},r_{c})=\mathbf{e}_{r^{\prime}}^{T}\sum_{r_{c}}\left(f(z^{i}_{gc})f(z^{i}_{cg})^{T}\right)\mathbf{e}_{r_{c}}.$
In summary, the entanglement between latent variables emerges from projecting onto a shared basis.
Next we turn to investigating entanglement in real-time. Since SII is an accumulating measure (best used for comparing full trajectories) it is not well suited for high temporal fidelity analysis. Instead, we aim to compare the Naive marginal likelihood $p_{Z}(\mathbf{o}_{t}\mid r_{g})$ and the history-conditioned marginal likelihood $p_{Z}(\mathbf{o}_{t}\mid r_{g},\mathbf{o}_{1:t-1})$. The history-conditioned marginal likelihood can be approximated empirically using the dynamics of the Optimal Observer’s marginalized beliefs:
$p_{Z}(\mathbf{o}_{t}\mid r_{g},\mathbf{o}_{1:t-1})=\sum_{r_{c}}p_{Z}(\mathbf{o}_{t}\mid r_{g},r_{c})p(r_{c}\mid r_{g},\mathbf{o}_{1:t-1})\approx\frac{B_{t+1}(r_{g})}{B_{t}(r_{g})}.$
Here we point out that the Optimal Bayesian update is equivalent to the Naive update ($\frac{B_{t+1}(r_{g})}{B_{t}(r_{g})}\propto p_{Z}(\mathbf{o}_{t}\mid r_{g})$) when the belief dynamics are Markovian with respect to the marginals, meaning that the transition from $B_{t}(r_{g})$ to $B_{t+1}(r_{g})$ is independent of $\mathbf{o}_{1:t-1}$.
The history-dependence of the Optimal Bayesian update can be attributed to the prior’s impact on the functional likelihood, $\sum_{r_{c}}p_{Z}(\mathbf{o}_{t}\mid r_{g},r_{c})p(r_{c}\mid r_{g},\mathbf{o}_{1:t-1})$, manifesting as entanglement between variables $r_{g}$ and $r_{c}$. To quantify this ”non-Markovian-ness”, we measure the Jeffreys Divergence, $J(p,q)={D}_{\mathrm{KL}}(p\;\|\;q)+{D}_{\mathrm{KL}}(q\;\|\;p)$, between (normalized) Naive marginal likelihoods and history-conditioned marginal likelihoods, which we refer to as dis-entanglement:
$\text{Dis-entanglement}_{t}=J\bigl(\frac{p_{Z}(\mathbf{o}_{t}\mid r_{g})}{z_{a}}\;||\;\frac{p_{Z}(\mathbf{o}_{t}\mid r_{g},\mathbf{o}_{1:t-1})}{z_{b}}\bigr)\approx\!J\bigl(\frac{p_{Z}(\mathbf{o}_{t}\mid r_{g})}{z_{a}}\;||\;\frac{B_{t+1}(r_{g})}{B_{t}(r_{g})z_{b}}\bigr)\!,$
where $z_{a}$ and $z_{b}$ are normalization factors to ensure valid probabilities.
To investigate the response of the network dynamics to dis-entanglement events, we performed a cross correlation between dis-entanglement of Optimal Bayesian beliefs and several properties of the network output.Figure 14a demonstrates that dis-entanglement is associated with a change of activity in population space, in both Fully Trained and Echo State networks. However, the magnitude of the response differs greatly between networks (Figure 14b). Dis-entanglement is associated with an increase in activity in the Fully Trained network, in contrast to Echo State activity which appears to shrink. Furthermore, Figure 14c demonstrates that the Fully Trained network’s increase in activity is distributed throughout the population, rather than along a few directions. The combined effect can be observed in Figure 14c, which shows the evolving Participation Ratio of the recurrent activity, which exhibits a spike in dimensionality for the Fully Trained network. This dynamic spread of information could account for the increased RNN dimensionality observed in Figure 11d. These results imply that active control of dimensionality may be a mechanism for disentanglement.
Figure 14: Dis-entanglement correlates with learned dynamics of dimensionality. Cross-correlations between Bayesian dis-entanglement and network (a) absolute distance, (b) L2 norm, (c) Hoyer’s sparsity and (d) Participation Ratio of network dynamics. Only data from $20\leq t\leq T$ was analyzed to isolate the steady-state response profile.
A.5 Representation Classification Chains and Thousand Brains Theory
The Thousand Brains Theory of intelligence [33] postulates that individual cortical columns autonomously model through compositions of compositions. Tasked specifically with learning useful relationships, Representation Classification Chains can be viewed as the potential microcircuit of such a system. This architecture-agnostic circuit generates abstractions $\mathbf{r}$, to explain inputs $\mathbf{o}$, in a manner that enables optimal control of inputs given their internal value, without needing to interact with the environment. These chains incorporate the relational information of embeddings, decision-making of Classifiers, latent representation of generative models and embodiment of a Controller. Importantly, these modules are not combined arbitrarily, but have a principled grounding in the task of compositional generalization to novel combinations of latent variables.
Figure 15: Additional learning trajectories of the Controller trained Offline w/ Generator. Representative examples of the Controller exploring an internally generated Cognitive Gridworld.
Figure 16: Network results extend to environments with 3 interacting variables. (a) Same as Figure 3b-d, for $1,2$ and 33 relevant variables. (b) Divergence of network marginal beliefs from Bayesian marginal beliefs. Over the course of inference, the Fully Trained network’s beliefs (solid lines) diverge from Naive Bayes (on the x-axis) while staying aligned with Joint Bayes (on the y-axis). Conversely, the Echo State network’s beliefs (dashed lines) exhibit the opposite trend, aligning closer to Naive Bayes beliefs. Line colors correspond to contextual variables $C=2$ and $C=3$.
A.6 Environment Hyperparameters
  • $T$ (Trajectory / inference steps): 30
  • $d_{\mathcal{E}}$ (Embedding dimensionality): 30
  • $|\mathcal{S}|$ (Total latent variables / states): 500
  • $R$ (Possible realizations): 10
  • $d_{o}$ (Observation dimensions): 5
  • $\lambda$ (Likelihood temperature): 2
A.7 Additional Model Hyperparameters
All models are implemented in PyTorch.
General.
  • Hidden dimensionality of all networks ($N$): 1000
  • Parametric embedding dimensionality ($d_{\hat{\mathcal{E}}}$): 1000
  • Optimizer: Adam
Experiment 1.
  • Learning rate: 0.001
  • Training episodes: 50000
  • Training batch size: 8000
Experiment 2.
  • RCC Learning rate: 0.0005
  • RCC Training episodes: 125000
  • RCC Training batch size: 8000
  • RCC Entropy bonus ($\beta_{1}$): $0.1$
Experiment 3.
  • RCC Learning rate: 0.0001
  • RCC training episodes: 50000
  • RCC Training batch size: 20000
  • RCC Entropy bonus ($\beta_{1}$): $0.01$
  • Controller Learning rate: 0.005
  • Controller Training episodes: 2000
  • Controller Batch size: 20000
  • Controller Entropy bonus ($\beta_{2}$): $0.05$
  • Online Generator Learning Rate: 0.001
Learnable Compositional Embeddings.
When compositional embeddings are learned, training reliability was improved when the learnable embedding estimates were over-parameterized. The relevant variable embedding estimates were passed through learnable linear layers $f_{K}(\cdot),f_{Q}(\cdot)\in\mathbb{R}^{d_{\hat{\mathcal{E}}}\times d_{\mathcal{E}}}$, to project them down from $d_{\hat{\mathcal{E}}}=N$ to dimensionality $d_{\mathcal{E}}$. Estimated interactions were computed from the projections:
$\hat{\mathbf{z}}^{i}=\langle f_{K}(\hat{K}^{i}),f_{Q}(\hat{Q}^{i})\rangle$
Embedding regularization was applied to the projections:
$\mathcal{L}_{\text{reg}}=\frac{1}{C\times d_{o}}\sum_{c,i}(\|f_{K}(\hat{\mathbf{k}}^{i}_{c})\|_{2}-1)^{2}+(\|f_{Q}(\hat{\mathbf{q}}_{c}^{i})||_{2}-1)^{2}$
Classifier Architecture.
  1. 1.
    Input: At each time step $t$, the input $x_{t}^{clss}$ is the concatenation of the current observation vector $o_{t}$ and the estimated interactions $Z$: $x_{t}^{cls}=[o_{t},Z]$.
  2. 2.
    Read-in: $h_{t}^{in}=\text{ReLU}(W_{in}x_{t}^{cls}+b_{in})$ maps the input to the hidden dimension $N$.
  3. 3.
    LSTM: $h_{t}^{stm},h_{t}^{ltm}=\text{LSTM}(h_{t}^{in},(h_{t-1}^{stm},h_{t-1}^{ltm}))$ processes the $T\times N$ sequence where initial hidden states $(h_{0}^{stm},h_{0}^{ltm})$ are learnable parameters.
  4. 4.
    Read-out: $l_{t}$: $l_{t}=W_{out}h_{t}^{stm}+b_{out}$ maps the LSTM output to logits of dimension $R$.
  5. 5.
    Output: As described in the main text, a cumulative sum is applied along time and then a softmax is applied along the realizations to produce the Classifier’s belief-state.
Generator Architecture.
To provide the Generator estimates of the relevant variable realizations we sample $\mathbf{\hat{r}}$ from the Classifier’s final belief-state $B_{T}$. We assume that the Classifier has already acted and insert the taken action $a$ into the $g$’th index of $\hat{\mathbf{r}}$. We also provide the Classifier’s confidence in the sampled realizations $\hat{\mathbf{r}}_{conf}=B_{T\hat{\mathbf{r}}}$. Depending on whether $a$ was rewarded or not, the $g$’th index of $\hat{\mathbf{r}}_{conf}$ is set to $1$ or 0.
Input: The Generator integrates three inputs: $x^{gen}=h_{r}+h_{c}+h_{z},\quad\text{where }h_{r},h_{c},h_{z}\in\mathbb{R}^{N}.$
[…]
  • $h_{c}=W_{c}\hat{\mathbf{r}}_{conf}+b_{c}$ is a linear projection of the confidence in the sampled realizations.
  • $h_{z}=W_{z}\hat{\mathbf{z}}^{i}+b_{z}$ is a linear projection of estimated interactions $\hat{\mathbf{z}}^{i}$. Since dimensions of the observation space are independent we process them in parallel.
Output: $\hat{P}_{\hat{Z}}(\mathbf{o}|\hat{\mathbf{r}})=\sigma\biggl(W_{3}\bigl(\text{ReLU}\bigr(W_{2}(\text{ReLU}(W_{1}x^{gen}+b_{1}))+b_{2}\bigr)\bigr)+b_{3}\biggr)$.
Training reliability was improved by augmenting the Generator’s loss function to give greater weight to elements of a batch where the Classifier’s action was rewarded. The augmented loss was:
$\mathcal{L}^{\star}_{\text{Generator}}=\text{chance}\times\mathcal{L}_{\text{Generator}}^{1-\delta}+\frac{1-\text{chance}}{\langle\delta(a-r_{g})\rangle}\times\mathcal{L}_{\text{Generator}}^{\delta}+\mathcal{L}_{\text{reg}},$
where $\text{chance}=\frac{1}{R}$ and $\mathcal{L}_{\text{Generator}}^{\delta}$ & $\mathcal{L}_{\text{Generator}}^{1-\delta}$ are the $\mathcal{D}_{\mathrm{KL}}$ losses where $a$ was rewarded and unrewarded, respectively, and $\langle\delta(a-r_{g})\rangle$ is the average accuracy in the batch.
Controller Architecture.
  1. 1.
    Input: The Actor and Critic networks pass $\hat{Z}$ through separate two-layer MLPs with ReLU activations. For the Actor, the final output is passed through a Softmax to produce a distribution over actions. For the Critic, the output is passed through a Sigmoid function to output the predicted intrinsic reward.
A.8 Pseudo-code
Algorithm 1 Cognitive Gridworld Task
Input: Total variables / states $|\mathcal{S}|$, Context size $C$, Realizations $R$, Time Horizon $T$
Hyperparameters: Temperature $\lambda$
Initialize Embeddings: $K,Q\leftarrow$ Create Compositional Embeddings($|\mathcal{S}|$)
Create Compositional Embeddings:
$K,Q\sim\mathcal{N}(0,I)$ {Draw matrix of size $|\mathcal{S}|\times d_{o}\times d_{\mathcal{E}}$}
for $s\in\{1,\dots,|\mathcal{S}|\}$ do
  $K_{s}\leftarrow\textsc{GramSchmidt}(K_{s})$ {Orthogonalize across $d_{o}$}
  $Q_{s}\leftarrow\textsc{GramSchmidt}(Q_{s})$
end for
[…]
return $K,Q$
Compression & Expansion(K,Q):
 Initialize sinusoid $v$ of size $R$ with wavelength $2R$ {Sinusoidal base}
for relevant variable pair $c,c^{\prime}\in C$ do
  $z\leftarrow\langle\mathbf{k}_{c},\mathbf{q}_{c^{\prime}}\rangle$ {Get cosine similarity}
  $\phi\leftarrow\arccos(z)$ {Convert sim to angle}
  $v_{cc^{\prime}}^{i}\leftarrow v(r+\phi)$ {roll base by angle}
end for
$\ell_{\mathbf{z}^{i}}\leftarrow\sigma\biggl(\sum\text{Broadcast}(\{v_{cc^{\prime}}^{i}(v_{c^{\prime}c}^{i})^{\top}\})\biggr)$ {Tensor of shape $R^{C}$ parameterized by ZZ}
return $\ell_{\mathbf{z}^{i}}$ {Likelihood of observing $^{\prime}0^{\prime}$ in dimension $o$ of $\mathbf{o}$}
Episode Generation:
 Sample context indices $C\subset\{1,\dots,|\mathcal{S}|\}$
 Sample realizations $r_{c}\sim\text{Uniform}(R)$ for $c\in C$
 Sample goal $g\sim\text{Uniform}(C)$
$a^{*}\leftarrow r_{g}$
for observation dimension $o\in\mathcal{O}$ do
  $\ell_{\mathbf{z}^{i}}\leftarrow\text{Compression \& Expansion}(K^{i},Q^{i})$
end for
 Generate trajectory $\mathbf{o}_{1:T}$ where $o_{t}^{i}\sim\text{Bernoulli}(\ell_{\mathbf{z}^{i}}(\mathbf{r}))$
Algorithm 2 Experiment 1: Training to capture Semantic Interaction Information
Input: Episodes $E$, Interactions $Z$, FullyTrainable or EchoState Network $\Phi$
for each episode $(\mathbf{o}_{1:T},r_{g},Z)\in E$ do
  $h_{1}\leftarrow h_{\phi}$ {Initialize memory streams}
  $\mathcal{L}\leftarrow 0$ {Initialize loss}
  $\ln B^{\text{Net}}_{1}\leftarrow\mathbf{0}$ {Initialize log belief}
  $\ln B^{\text{Joint}}_{1}\leftarrow\mathbf{0}$ {Initialize log belief}
  for $t=1$ to T1T-1 do
   $h_{t+1},\Delta]\leftarrow\Phi([\mathbf{o}_{t},Z],h_{t})$
   $\ln B_{t+1}^{\text{Net}}\leftarrow\ln B_{t}^{\text{Net}}+\Delta$
   $\ln B_{t+1}^{\mathrm{Joint}}\leftarrow\ln B_{t}^{\mathrm{Joint}}+\ln P_{Z}(\mathbf{o}_{t}|r_{g},r_{c})-\ln P_{Z}(\mathbf{o}_{t})$
   $\mathcal{L}\leftarrow\mathcal{L}+\frac{1}{T}\mathcal{D}_{KL}\bigl(\text{Softmax}(B_{t+1}^{\text{Joint}})||\text{Softmax}(B_{t+1}^{\text{Net}})\bigr)$
  end for
  Update $\Phi,h_{\phi}$ to minimize $\mathcal{L}$
end for
Algorithm 3 Experiment 2: Training Representation Classification Chains
Input: Episodes $E$, Classifier $\Phi$, Generator $\Psi$
Hyperparameters: Entropy bonus weight $\beta$
Learnable Embeddings: $\hat{K},\hat{Q}$, $f_{K},f_{Q}$
for each episode $(\mathbf{o}_{1:T},r_{g},C)\in E$ do
  $\hat{Z}\leftarrow\langle f_{K}(\hat{K}_{C}),f_{Q}(\hat{Q}_{C})\rangle$ {Estimated Interactions}
  $\ln B^{\text{Net}}_{1}\leftarrow\mathbf{0}$
  $h_{1}\leftarrow h_{\phi}$
  $\Phi$ Forward Pass:
  for $t=1$ to T1T-1 do
   $h_{t+1},\Delta]\leftarrow\Phi([\mathbf{o}_{t},\text{Detach}(\hat{Z})],h_{t})$
   $\ln B_{t+1}^{\text{Net}}\leftarrow\ln B_{t}^{\text{Net}}+\Delta$
  end for
  Sample action $a\sim B_{Tg}^{\text{Net}}$ {Action taken at t=Tt=T}
  $\delta\leftarrow\mathbb{I}(a=r_{g})$ {Sparse reward signal}
  $\Phi$ Backward Pass:
  $\mathcal{L}_{Clss}\leftarrow 0$
  for $t=2$ to $T$ do
   $\text{EB}\leftarrow-B_{tga}^{\text{Net}}\ln B_{tga}^{\text{Net}}$ {Entropy Bonus}
   $\text{PG}\leftarrow\delta\ln B_{tga}^{\text{Net}}+(1-\delta)\ln(1-B_{tga}^{\text{Net}})$ {Policy Gradient}
   $\mathcal{L}_{Clss}\leftarrow\mathcal{L}_{Clss}-\frac{1}{T}(\text{PG}+\beta\cdot\text{EB})$
  end for
  Update $\Phi,h_{\phi}$ to minimize $\mathcal{L}_{Clss}$ {Gradient flows through $\Phi$}
  $\Psi$ Forward Pass:
  $\hat{r}_{g}\leftarrow a$
  $\hat{r}_{c}\sim B_{Tc}^{\text{Net}}$ {Sample context realization}
  $\hat{P}_{\hat{Z}}(\mathbf{o}\mid\hat{\mathbf{r}})\leftarrow\Psi([\hat{Z},\hat{r}_{g},\hat{r}_{c}])$
  $\Psi$ Backward Pass:
  $\mathbf{OPE}^{i}\leftarrow\mathcal{D}_{\text{KL}}(\hat{P}_{\hat{Z}}(o\mid\hat{\mathbf{r}}),\tilde{P}(\mathbf{o}))$ {Observation Prediction Error}
  $\mathbf{REG}^{i}_{V}\leftarrow\langle\text{MSE}(\|f_{V}(V^{i})\|_{2},1)\rangle_{c}$ {Embedding Regularization}
  $\mathcal{L}_{Gen}\leftarrow\langle\mathbf{OPE}+\mathbf{REG}_{\hat{K}}+\mathbf{REG}_{\hat{Q}}\rangle_{o}$
  Update $\Psi,\hat{K}_{C},\hat{Q}_{C},f_{K},f_{Q}$ to minimize $\mathcal{L}_{Gen}$ {Gradient flows through $\Psi$ to Embeddings}
end for
Algorithm 4 Experiment 2: Offline Controller Training
Input: Learned Factors $\hat{Z}$, Generator $\Psi$, Preferred Observation $\Omega$
Hyperparameters: Training Episodes $E$, Entropy bonus weight $\beta_{cntrl}$, Bonus decay $\beta_{\text{decay}}$
Models: Actor $\Pi$, Critic $V$
Intrinsic Reward Function $\mathcal{R}(\mathbf{r},\mathbf{\Omega},\Psi,Z)$:
$\ell_{\mathbf{z}^{i}}\leftarrow\Psi(\hat{\mathbf{z}}^{i},\mathbf{r})$ {Generator provides likelihood estimates}
$\text{PL}^{i}=\ln(\ell_{\mathbf{z}^{i}})\Omega^{i}+\ln(1-\ell_{\mathbf{z}^{i}})(1-\Omega^{i})$ {Preference logit}
$\text{Reward}\leftarrow\exp\langle\text{PL}\rangle_{o}$
return Reward
for $e=1$ to $E$ do
[…]
  $\textbf{PE}\leftarrow-\sum_{\mathbf{r}}\pi(\mathbf{r})\ln\pi(\mathbf{r})$ {Compute policy entropy}
  $\mathbf{r}\sim\pi(\mathbf{r})$ {Sample joint action}
[…]
  $\mathcal{L}_{cntrl}\leftarrow A^{2}-\ln\pi(\mathbf{r})\cdot\text{Detach}(A)-\beta_{cntrl}\cdot\textbf{PE}$
  $\beta_{cntrl}\leftarrow\beta_{cntrl}\cdot\beta_{\text{decay}}$ {Decay entropy bonus}
  Update $\Pi,V$ to minimize $\mathcal{L}_{cntrl}$
end for
Statement on LLM usage
Large language models (LLMs) were used to assist with improving the clarity and style of the manuscript text including this statement; Additionally, LLMs aided in providing relevant articles. All conceptual contributions, experimental design, and analysis were conducted by the author.